OpenStudy (anonymous):

A porch has the shape of a quarter circle with a diameter of 5 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate for Pi(3.14) .

OpenStudy (igreen):

Find the area: \(A = \pi r^2\)

OpenStudy (anonymous):

Then divide by 4

OpenStudy (anonymous):

Since it's a quarter circle

OpenStudy (igreen):

The diameter is 5, so the radius is 2.5 \(A = (3.14)(2.5^2)\)

OpenStudy (igreen):

Or: \(A = \dfrac{(3.14)(2.5^2)}{4}\)

OpenStudy (anonymous):

4.90625

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (anonymous):

Thanks to the both of you

OpenStudy (anonymous):

No problem. :)